OpenStudy (anonymous):

Simplify the following expression. (58x + 64y) - (8x - 21y) 50x + 43y 50x - 43y 50x - 85y 50x + 85y

OpenStudy (anonymous):

What do you think ? We can sat that : \[(58x+64y)-(8x-21y)=58x+64y-8x+21y=(58x-8x)+(64y-21y)\] Can you complete ?

OpenStudy (anonymous):

c.?

OpenStudy (anonymous):

50x - 85y

OpenStudy (anonymous):

NO ! You have : \[64y-21y\] What does it equal ?

OpenStudy (cicilybailey14):

Let's simplify step-by-step. 58x+64y−(8x−21y) Distribute the Negative Sign: =58x+64y−(8x−21y) =58x+64y+−8x+21y =58x+64y+−8x+21y Combine Like Terms: =58x+64y+−8x+21y =(58x+−8x)+(64y+21y) =50x+85y Answer: =50x+85y

OpenStudy (cicilybailey14):

:)

OpenStudy (anonymous):

D my bad i was looking at the wrong one but i ment d i saw the word problem and was oohh...

OpenStudy (cicilybailey14):

haha :D